OpenStudy (anonymous):

Evaluate cos(cos^-1 3/5 + sin^-1 5/13)

OpenStudy (anonymous):

=(12/13)(3/5)-(5/13)(4/5) =36/65-20/65 =16/65 can someone verify, thx

OpenStudy (ranga):

Yes, I am getting 16/65 too.

OpenStudy (anonymous):

awesome, thx

OpenStudy (ranga):

you are welcome.